The magnetic moment of an electron revolving in nth orbit is directly proportional to

Grade:12

1 Answers

Khimraj
6 years ago
Let the electron is revolving in in orbit of radius r and with velocity v.
magnetic moment of elecron is given by
μ = iA
i = q/T = -e/T
T = 2πr/v
So i = -ev/2πr
A = πr2
So μ = iA = -evr/2
μ = -evrme/2me = (-e/2me)*L , where L = angular momentum = mvr
Angular momentum of electron revolving in atom is given by
L = nh/2π
μ So  μ = n(eh/4πme).
So magnetic moment of electron revolving in nth orbit is directly proportional to n.
Hope it clears. If you like answer then please approve it.